OpenStudy (anonymous):

The width of a rectangular is 6 kilometers less than twice its length. If its area is 216 square kilometers, find the dimensions of the rectangle.

OpenStudy (anonymous):

w=2L-6 Since A=w*L A=(2L-6)*L We know A=216, therefore L*(2L-6)=216 Solve this and you get the length, and use the first equation to get the width.